OpenStudy (anonymous):

If sin (A) = 5/8 with A in QII, find sec (2A).

OpenStudy (gorv):

sec(2A)=1/cos(2A)

OpenStudy (gorv):

and cos(2A)=1-2sin^2(A)

OpenStudy (gorv):

can u do it now???

OpenStudy (anonymous):

actually...i worked on it I just didnt get the right problem

OpenStudy (anonymous):

I got \[\frac{ 8\sqrt{2} }{ 5}\]

OpenStudy (gorv):

follow me

OpenStudy (gorv):

cos(2A)=1-2*(sinA)^2

OpenStudy (gorv):

sinA=5/8 (sinA)^2=??

OpenStudy (anonymous):

ok let me work it out...

OpenStudy (gorv):

okk

OpenStudy (anonymous):

7/32?